Current Position Analysis

The following data were taken from the balance sheet of Nilo Company at the end of two recent fiscal years:

Current Year Previous Year
Current assets:
  Cash $433,200 $353,600
  Marketable securities 501,600 397,800
  Accounts and notes receivable (net) 205,200 132,600
  Inventories 752,400 444,100
  Prepaid expenses 387,600 283,900
  Total current assets $2,280,000 $1,612,000
Current liabilities:
  Accounts and notes payable
  (short-term) $348,000 $364,000
  Accrued liabilities 252,000 156,000
  Total current liabilities $600,000 $520,000

a. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round ratios to one decimal place.

Current Year Previous Year
1. Working capital $ $
2. Current ratio
3. Quick ratio

b. The liquidity of Nilo has from the preceding year to the current year. The working capital, current ratio, and quick ratio have all . Most of these changes are the result of an in current assets relative to current liabilities.
Answer #1

(a) Working capital = Current asset- Current liabilities

So,for current year Working capital = 22,80,000- 6,00,000=16,80,000

For previous year working capital=16,12,000-5,20,000=10,92,000

(b) Current ratio= Current asset/Current liabilities

Second part of question is not much clear. Ask it with clarity

So for current year Current ratio= 22,80,000/6,00,000=3.8 or 4 approx

for previous year= 16,12,000/5,20,000=3.1 or 3 approx

(C) Quick ratio=Quick asset/Current liabilities

Quick assets= Current asset- Inventories-Prepaid stock

For current year( quick assets)= 22,80,000-7,52,400-3,87,600=11,40,000

So quick ratio=11,40,000/600000=1.9 or 2 approx

For previous year (quick assets)=16,12,000-4,44,100-2,83,900=8,84,000

Quick ratio=8,84,000/5,20,000=1.7 or 2 approx